OpenStudy (anonymous):

3x – 9y = -18 –x + 3y = 6 need to solve for x and y

sam (.sam.):

3x – 9y = -18................x(-1) –x + 3y = 6....................x(3) ------------------------------------------ -3x+9y = 18 -3x+9y = 18 --------------------------------------- No solution